I am totally clueless! please help me figure out this word problem. If the height of a right circular cylinder is quadrupled and the radius is divided by three, how is the volume changed?
Answer by ptaylor(2198)


Volume of a right circular cylinder is given by the formula:
V=pi*r^2*h where r is radius and h is height
Volume of original cylinger is: V1=pi*r^2*h
Volume of new cylinder is: V2=pi*(r/3)^2*(4h); simplifying, we get:
V2=pi*((r^2)/9)*4h=
V2=(pi*r^2*h)*(4/9) but (pi*r^2*h)=V1, so we have:
V2=(4/9)*V1
So the Volume gets quite a bit smaller---4/9 of the original volume

Hope this helps---ptaylor